At the beginning of the period, the Cutting Department budgeted direct labor of $125,000, direct materials of $167,000 and fixed factory overhead of $12,000 for 7,800 hours of production. The department actually completed 11,200 hours of production. What is the appropriate total budget for the department, assuming it uses flexible budgeting? Round hourly rates to two decimal places.

Round interim calculations to two decimal places. Round your final answer to the nearest dollar.

a.$304,000

b.$436,513

c.$309,231

d.$431,328
Answer #1
Direct labor per hour = Direct labor/budgeted hour
Direct labor per hour = 125000/7800 hours = 16.02
Direct material per hour = 167000/7800 hours = 21.41
Fixed factory hour = 12000
Calculation of total budget for department
direct labor (11200*16.02) 179424
Direct material (11200*21.41) 239792
Fixed factory overhead 12000
total budget 431216
it is nearest to option "d" 431328